OpenStudy (anonymous):

what is the slope of the line which passes through (-6,0) and (4,0)

OpenStudy (anonymous):

\[m= \frac{y1-y2}{x2-x2} \] \[ m=\frac{0-0}{4+6} \] m=0/10 m=0 Slope = 0

OpenStudy (amistre64):

ohhh, its a harry potter themed icon.

OpenStudy (anonymous):

Yes, it is :)